OpenStudy (anonymous):

John is going to remodel his bed room. He is going to put in new carpet, new trim, and he is going to paint the walls. The dimensions of his room are 9.5'nby 12', and the room has an 8 foot ceiling. The cost to do the carpet is $7.35 per square foot. The trim will cost $4.90 per linear foot, and the cost to paint the walls will be $2,61 per square foot/ What will be the total cost of his project?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

step#1 we have to compute the perimeter of the room: thatperimeter is: 2*(12+9.5)=...?

OpenStudy (anonymous):

43

OpenStudy (michele_laino):

ok! so the cost for trim is: 43*4.90=...?

OpenStudy (anonymous):

2107

OpenStudy (michele_laino):

I got $210.7

OpenStudy (anonymous):

me to

OpenStudy (anonymous):

i forgot to add decimal

OpenStudy (michele_laino):

step#2 we have to compute the area of tha room floor. That area is: A=12*9.5=...?

OpenStudy (anonymous):

114

OpenStudy (michele_laino):

so the cost of carpet, is: 114*7.35=...?

OpenStudy (anonymous):

837.9

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

step#3 now we have to compute the area of the surface of the room. That area is: A=2*(12*9.5+9.5*8+12*8)=...?

OpenStudy (anonymous):

371.5

OpenStudy (michele_laino):

I got 572

OpenStudy (anonymous):

your prolly right

OpenStudy (anonymous):

yep i got 572 after imputing it again

OpenStudy (michele_laino):

so the cost for paint, is: 572*2.61=...?

OpenStudy (anonymous):

1492.92

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

then, the final cost, is: 1492.92+837.9+2107=...?

OpenStudy (michele_laino):

please wait, I have made an error

OpenStudy (michele_laino):

he last area is not 572, it is: A=2*(9.5*8+12*8)+12*9.5=...?